# SARCLIP: A Vision Language Foundation Model for Semantic Understanding and Target Recognition in SAR Imagery

### v1

@@ III-C Model Training
where $\text{sim}(a,b)=\frac{a^{\top}b}{\|a\|\|b\|}$ is cosine similarity, $\tau$ donates a temperature hyperparameter, and $N$ is the batch size.
III-C2 Downstream fine-tuning.
@@
TABLE III: Retrieval performance on the SARCLIP-1M test set and improvements over the SARCLIP baseline(%). LAION is an optical image dataset in the natural scene domain, while RS5M, Skyscript, and R3+D10+S4 are optical image datasets in the remote sensing domain. , §, and denote models pretrained on RS5M, Skyscript, and R3+D10+S4, respectively.
@@
SARCLIP ResNet-50 LAION SARCLIP-1M 2.71 10.82 17.57 3.07 11.76 19.03 10.49
SARCLIP ResNet-50 R3+D10+S4 SARCLIP-1M 2.98 10.79 17.35 3.21 12.02 19.39 10.95 (+4.38)
HarMA ViT-B-32 LAION SARCLIP-1M 1.15 4.80 9.00 1.13 5.05 9.21 5.06
SARCLIP ViT-B-32 LAION SARCLIP-1M 3.42 12.82 20.32 3.63 13.72 21.40 12.55
SARCLIP ViT-B-32 RS5M SARCLIP-1M 3.73 13.33 20.61 3.95 13.99 21.77 12.89
SARCLIP § ViT-B-32 SkyScript SARCLIP-1M 3.79 13.23 20.31 3.72 13.80 21.46 12.72
SARCLIP ViT-B-32 R3+D10+S4 SARCLIP-1M 3.58 13.25 20.89 3.79 13.64 21.52 12.77 (+1.75)
SARCLIP ViT-L-14 LAION SARCLIP-1M 4.73 15.53 23.81 4.75 16.09 24.51 14.90
SARCLIP ViT-L-14 RS5M SARCLIP-1M 4.81 15.60 23.57 4.77 16.56 25.26 15.09
SARCLIP § ViT-L-14 SkyScript SARCLIP-1M 4.69 15.70 23.84 4.67 16.34 24.65 14.98
SARCLIP ViT-L-14 R3+D10+S4 SARCLIP-1M 4.69 15.87 24.04 4.90 16.62 25.42 15.25 (+2.34)
IV Experiment
IV-A Datasets
To evaluate the superiority of our proposed SARCLIP, three extensively adopted SAR target recognition datasets and our SARCLIP-1M test retrieval dataset is utilized for downstream performance evaluation.
IV-A1 SARCLIP-1M test.
This SAR retrieval dataset comprises 20,628 image-text pairs compiled from existing datasets. The test set consists of 180 images from MSTAR, 6,667 images from ATRNet-STAR, 2,783 images from SARDet-100K, 7,096 images from FAIR-CSAR, and 3,902 images from SARLANG-1M-Captions. Each image is uniquely captioned.
IV-A2 MSTAR-SOC.
This SAR target recognition dataset is acquired by an X-band radar operating in HH polarization mode with a resolution of 0.3 meters. It comprises 10 military vehicle targets and is partitioned into 4 experimental settings, as established by [40]. Specifically, The SOC setting has 2,747 training images (17 angles) and 2,425 test images (15 angles), with all 10 classes present in both sets.
IV-A3 FUSAR-ship.
It is a dataset designed for ship recognition, comprising 15 main ship classes, 98 subclasses, and various non-ship maritime targets [15]. It is built from 126 Gaofen-3 images captured in ultrafine resolution mode (1.124 $\times$ 1.728 m) with dual-polarization (DH and DV). The dataset covers a wide range of scenarios, including open sea, coastal areas, rivers, islands, and land backgrounds.
IV-A4 SAR-VSA.
This SAR target recognition dataset comprises 25 fine-grained categories, integrating data from MSTAR, FUSAR-ship , and SAR-ACD [41]. It contains 11,045 training images and 8,161 testing images, as constructed by [15].
IV-B Evaluation Metrics
@@
where $\hat{y}_{i}$ is the predicted result and $y_{i}$ is the ground-truth label.
TABLE IV: Recognition results on MSTAR-SOC and SAR-VSA dataset and improvements over the OpenCLIP(%)($\mathbf{\%}$).
@@
IV-C Implementation Details
We develop SARCLIP based on the OpenCLIP framework. Automatic mixed-precision (AMP) training is employed to reduce memory usage. We adopt ResNet-50, ViT-B-32, and ViT-L-14 as image backbones, with learning rates set to 5e-4, 5e-5, and 5e-5, respectively. ResNet-50 is trained for 30 epochs, while ViT-B-32 and ViT-L-14 are trained for 10 epochs. The batch size is set to 256. Training is accelerated using the Adam optimizer, combined with a linear warm-up and cosine learning rate schedule. Downstream experiments are conducted using the wise-ft [42] framework. For the downstream target recognition task, we freeze the image backbone and fine-tune only the linear classification layers for 10,000 epochs. All experiments are conducted on two 80GB NVIDIA H100 GPUs.
IV-D Comparing with other Methods
In this section, we compare out method with other state-of-the-art CLP-based method include OpenCLIP, RemoteCLIP, GeoRSCLIP, SkyCLIP and HarMA [43].
TABLE V: Zero-shot classification results on FUSAR-ship and improvements over the RemoteCLIP baseline $(\mathbf{\%})$.
@@ IV-D Comparing with other Methods
IV-D1 Retrieval results on SARCLIP-1M test dataset.
@@
As presented in Table III, we compare our SARCLIP method with several state-of-the-art VLFMs designed for remote sensing. The results clearly indicate that existing models exhibit poor performance in SAR image-text retrieval tasks. In contrast to standard CLIP models (e.g., OpenCLIP), our training strategy effectively transfers knowledge from the optical domain (e.g. RS5M) to the SAR domain (SARCLIP-1M). Specifically, SARCLIP achieves a 15.25% mean recall while SARCLIP at 14.90%, demonstrating a significant improvement and highlighting the efficacy of our approach in multi-modal alignment in SAR domain.
@@
TABLE VI: Ablation study result on several downstream task. Img. Enc. and Text Enc. donate the image encoder and text decoder respectively.(%)($\mathbf{\%}$)
Model Architecture SARCLIP-1M test SOC VSA FUSAR
@@
IV-D2 Target recognition results on MSTAR-SOC and SAR-VSA dataset.
Table IV presents the performance of our SARCLIP models on downstream target recognition tasks after fine-tuning a linear layer, thereby demonstrating their visual understanding capabilities. The experimental results show that our SARCLIP series models consistently achieve superior performance. On the MSTAR-SOC dataset, SARCLIP achieves an impressive 99.83% accuracy, significantly outperforming optical VLFMs which typically hover around 80%. Similarly, on the SAR-VSA dataset, our model reaches 89.75% accuracy (specifically, SARCLIP with ViT-L-14), notably surpassing OpenCLIP’s 81.58%. These results underscore the strong visual feature extraction and understanding capabilities of our SARCLIP models for SAR imagery.
IV-D3 Zero-shot classification results on FUSAR-ship dataset.
Table V presents zero-shot classification results on the FUSAR-ship dataset, highlighting the generalization ability of SARCLIP without fine-tuning. While RemoteCLIP performs poorly (e.g., 1.19% for ViT-B-32), our SARCLIP models exhibit notably stronger zero-shot capabilities. In particular, SARCLIP with a ViT-L-14 backbone achieves the best performance at 12.41%, demonstrating improved transferability to unseen SAR ship categories. These results suggest that pre-training on a diverse SAR vision-language dataset like SARCLIP-1M significantly enhances zero-shot generalization over models trained primarily on visible-light data.
Figure 6: Ablation study on training layers of SARCLIP on the SARCLIP-1M test set (Mean Recall %).
IV-E Ablation Studies
IV-E1 Effect of the training encoder.
This section presents an ablation study evaluating the contributions of the image and text encoders to model performance across various downstream tasks. As summarized in Table VI, when only the text encoder is used, the model performs poorly, achieving just 5.23% mean recall on the SARCLIP-1M test set and 1.49% accuracy on FUSAR, highlighting the limited discriminative power of text alone for SAR imagery understanding. In contrast, activating only the image encoder significantly improves performance, with mean recall reaching 9.57%, and accuracy on MSTAR-SOC and SAR-VSA increasing to 99.62% and 88.94%, respectively demonstrating the image encoder’s effectiveness in SAR feature extracting. The combination of both encoders yields the best results, achieving 15.25% mean recall on SARCLIP-1M and 99.83% accuracy on SOC. Notably, zero-shot accuracy on FUSAR improves to 12.41%, confirming that cross-modal contrastive learning is essential for aligning semantic representations and enhancing generalization in both retrieval and zero-shot classification tasks.
IV-E2 Effect of the training layer.
As shown in Fig. 6, we evaluate the impact of varying the number of activated layers in the image and text encoders on the SARCLIP-1M test set. The results reveal a consistent trend: model performance improves as more layers are activated. This highlights the importance of deep, expressive representations in both modalities for effective SAR image-text understanding.
IV-F Visualization
As shown in Fig. 5, we utilize t-SNE [44] technique to visualize features extracted by SARCLIP (ViT-L-14) on three downstream datasets. The visualizations plots exhibit well-separated clusters across classes, demonstrating the model’s strong feature discrimination and providing insights into its visual and semantic representations.
V Conclusion
We present SARCLIP-1M, a large-scale SAR image-text dataset, and propose SARCLIP, the first vision-language foundation model for SAR. Trained via a two-stage domain transfer strategy, SARCLIP effectively transfers knowledge from optical remote sensing data to SAR imagery, achieving strong multi-modal alignment, target recognition and zero-shot capability. In the future, we will explore integrating multi-modal large language models and agent techniques for SAR image interpretation.

### v2

@@ 3.3.1 SARCLIP training
where $\text{sim}(a,b)=\frac{a^{\top}b}{\|a\|\|b\|}$ is cosine similarity, $\tau$ denotes a temperature hyperparameter, and $N$ is the batch size.
3.3.2 SARCap training
For the SAR image captioning task, we use the CoCa [Coca] architecture, which combines contrastive loss and captioning loss. The captioning loss minimizes the difference between the predicted and ground-truth captions:
$\mathcal{L}_{\text{caption}}=-\sum_{t=1}^{T}\log P(w_{t}|w_{1},...,w_{t-1},z_{v})$ (2)
where $T$ is the caption length, $w_{t}$ is the $t$-th token, and $P(\cdot)$ denotes the predicted probability of $w_{t}$ conditioned on the previous tokens and the visual feature.
3.3.3 Downstream fine-tuning
@@
Table 2: Retrieval performance on the SARVLM-1M test set and improvements over the SARCLIP baseline(%). LAION [OpenCLIP] is an optical image dataset in the natural scene domain, while RS5M, Skyscript, and R3+D10+S4 are optical image datasets in the remote sensing domain. , §, and denote models pretrained on RS5M [RS5M], Skyscript [Skyscript], and R3+D10+S4 [RemoteCLIP], respectively.
@@
SARCLIP ResNet-50 LAION SARVLM-1M 2.71 10.82 17.57 3.07 11.76 19.03 10.49
SARCLIP ResNet-50 R3+D10+S4 SARVLM-1M 2.98 10.79 17.35 3.21 12.02 19.39 10.95 (+4.38)
HarMA ViT-B-32 LAION SARVLM-1M 1.15 4.80 9.00 1.13 5.05 9.21 5.06
SARCLIP ViT-B-32 LAION SARVLM-1M 3.42 12.82 20.32 3.63 13.72 21.40 12.55
SARCLIP ViT-B-32 RS5M SARVLM-1M 3.73 13.33 20.61 3.95 13.99 21.77 12.89
SARCLIP § ViT-B-32 SkyScript SARVLM-1M 3.79 13.23 20.31 3.72 13.80 21.46 12.72
SARCLIP ViT-B-32 R3+D10+S4 SARVLM-1M 3.58 13.25 20.89 3.79 13.64 21.52 12.77 (+1.75)
SARCLIP ViT-L-14 LAION SARVLM-1M 4.73 15.53 23.81 4.75 16.09 24.51 14.90
SARCLIP ViT-L-14 RS5M SARVLM-1M 4.81 15.60 23.57 4.77 16.56 25.26 15.09
SARCLIP § ViT-L-14 SkyScript SARVLM-1M 4.69 15.70 23.84 4.67 16.34 24.65 14.98
SARCLIP ViT-L-14 R3+D10+S4 SARVLM-1M 4.69 15.87 24.04 4.90 16.62 25.42 15.25 (+2.34)
4 Experiment
4.1 Datasets
To evaluate the superiority of our proposed SARCLIP, we utilize downstream dataset as follows: SARVLM-1M test. This SAR retrieval dataset comprises 20,628 image-text pairs compiled from existing datasets. The test set consists of 180 images from MSTAR, 6,667 images from ATRNet-STAR, 2,783 images from SARDet-100K, 7,096 images from FAIR-CSAR, and 3,902 images from SARLANG-1M-Captions. Each image is uniquely captioned.
MSTAR-SOC. This SAR target recognition benchmark acquired by an X-band HH radar at 0.3 m resolution, containing 10 military vehicle classes and four standard experimental settings [AConvNet]. In the SOC setting, all 10 classes appear in 2,747 training images from 17 aspect angles and 2,425 test images from 15 angles.
FUSAR-Ship. This ship recognition dataset comprising 15 main ship classes, 98 subclasses, and diverse non-ship maritime targets [SARATR]. It is built from 126 Gaofen-3 ultrafine-resolution images (1.124 $\times$ 1.728 m, dual-pol DH/DV), covering open sea, coastal waters, rivers, islands, and land backgrounds.
SAR-VSA. This SAR target recognition dataset comprises 25 fine-grained categories, integrating data from MSTAR [MSTAR], FUSAR-ship [FUSAR-ship], and SAR-ACD [SAR-ACD]. It contains 11,045 training images and 8,161 testing images, as constructed by [SARATR].
AIR-SLT. This is a remote sensing semantic localization benchmark with 22 large-scale images and 59 text-query test cases, each paired with ground-truth bounding boxes [SeLo, SeLov2].
4.2 Evaluation Metrics
@@
where $\hat{y}_{i}$ is the predicted result and $y_{i}$ is the ground-truth label. For the semantic localization task, we adopt Rsu, Rda, Ras, and Rmi as evaluation metrics, following [SeLo]. For the image captioning task, we employ BLEU [bleu], ROUGE_L [rouge_L], METEOR [meteor], and CIDEr [cider].
4.3 Implementation Details
We develop SARCLIP based on the OpenCLIP framework. Automatic mixed-precision (AMP) training is employed to reduce memory usage. We adopt ResNet-50, ViT-B-32, and ViT-L-14 as image backbones, with learning rates set to 5e-4, 5e-5, and 5e-5, respectively. ResNet-50 is trained for 30 epochs, while ViT-B-32 and ViT-L-14 are trained for 10 epochs. The batch size is set to 256. Training is accelerated using the Adam optimizer, combined with a linear warm-up and cosine learning rate schedule. Downstream experiments are conducted using the wise-ft [wise-ft] and SLM[SeLo] framework. For the downstream target recognition task, we freeze the image backbone and fine-tune only the linear classification layers for 10,000 epochs. All experiments are conducted on two 80GB NVIDIA H100 GPUs. For SAR imagery caption task, we employ optical image caption dataset HQRS-210K [HQRS-210K] as pretrained data.
4.4 Comparing with other Methods
In this section, we compare our method with other state-of-the-art CLIP-based and CoCa-based method on downstream tasks.
Table 3: Recognition results on MSTAR-SOC and SAR-VSA dataset and improvements over the OpenCLIP(%)($\mathbf{\%}$).
@@ 4.4 Comparing with other Methods
Table 4: Zero-shot classification results on FUSAR-ship and improvements over the RemoteCLIP baseline $(\mathbf{\%})$.
@@
Table 5: Semantic Localization results on AIR-SLT datase.
[…]
RemoteCLIP ViT-L-14 0.7707 0.6722 0.2876 0.7257
SARCLIP ViT-L-14 0.5643 0.4830 0.5631 0.5000
SARCLIP ViT-L-14 0.5654 0.4531 0.5601 0.4889
SARCLIP § ViT-L-14 0.5675 0.4792 0.5677 0.4981
SARCLIP ViT-L-14 0.5741 0.4173 0.5731 0.4834
4.4.1 Retrieval results on SARVLM-1M test dataset
@@
As presented in Table 2, we compare our SARCLIP methods with several state-of-the-art VLFMs designed for remote sensing. The results clearly indicate that existing models exhibit poor performance in SAR image-text retrieval tasks. In contrast to standard CLIP models (e.g., OpenCLIP), our training strategy effectively transfers knowledge from the optical domain (e.g. RS5M) to the SAR domain (SARVLM-1M). Specifically, SARCLIP achieves a 15.25% mean recall while SARCLIP at 14.90%, demonstrating a significant improvement and highlighting the efficacy of our approach in multi-modal alignment in SAR domain.
4.4.2 Target recognition results on MSTAR-SOC and SAR-VSA dataset
Table 3 presents the performance of our SARCLIP models on downstream target recognition tasks after fine-tuning a linear layer, thereby demonstrating their visual understanding capabilities. The experimental results show that our SARCLIP series models consistently achieve superior performance. On the MSTAR-SOC dataset, SARCLIP achieves an impressive 99.83% accuracy, significantly outperforming optical VLFMs which typically hover around 80%. Similarly, on the SAR-VSA dataset, our SARCLIP reaches 89.75% accuracy, notably surpassing OpenCLIP’s 81.58%. These results underscore the strong visual feature extraction our SARCLIP models for SAR imagery.
Table 6: Image Caption results on SARVLM-1M test datset with image encoder(ViT-L-14)(%)($\mathbf{\%}$).
Method Image Encoder Pretrain Data Tune On BLUE-1 BLUE-2 BLUE-3 BLUE-4 METEOR ROUGE_L CIDEr
CoCa-Vanilla ViT-L-14 LAION 8.75 3.21 1.14 0.32 10.54 10.67 4.17
CoCa-HQRS ViT-L-14 LAION HQRS 10.79 4.02 1.14 0.39 8.21 13.11 7.47
SARCap ViT-L-14 SARVLM-1M 22.42 14.44 10.45 7.79 19.23 21.76 24.32
SARCap ViT-L-14 LAION+HQRS SARVLM-1M 26.10 17.13 11.93 8.47 22.18 26.15 29.58
4.4.3 Zero-shot classification on FUSAR-ship dataset
Table 4 summarizes zero-shot classification on FUSAR-ship, evaluating SARCLIP without fine-tuning. RemoteCLIP performs poorly (e.g., 1.19% for ViT-B-32), whereas SARCLIP variants show much stronger zero-shot recognition. In particular, SARCLIP with a ViT-L-14 backbone reaches 12.41%, indicating better transfer to unseen SAR ship categories. This demonstrates that pre-training on the large-scale SARVLM-1M SAR vision-language corpus greatly enhances zero-shot generalization over models trained mainly on visible-light data.
4.4.4 Semantic localization on AIR-SLT dataset
Table 5 reports semantic localization results on the optical AIR-SLT dataset under a cross-domain setting. Our SARCLIP achieves 0.5000 Rmi compared to 0.7257 for RemoteCLIP, showing that SARCLIP retains a certain level of generalization to optical scenes. As illustrated in Figure 7, given query captions such as “four white planes” and ”white airport”, our model successfully captures the plane features but misses the ”white airport” region, which is mainly attributed to the intrinsic differences between SAR and optical imagery.
4.4.5 SAR imagery caption on SARVLM-1M
Table 8 presents the caption result from our SARCap framework. After high-quality optical remote sensing fine-tuning, CoCa-HQRS performs better than CoCa-Vanilla. We utilize a progressive training strategy to obtain our SAR imagery captioning model SARCap which outperforms SARCap from scripts. As shown in Figure 8, our SARCap generate fluent and meaningful caption for different SAR sences.
4.5 Ablation Studies
4.5.1 Effect of downmain transferring strategy
There is a large perspective gap between natural and optical remote sensing images, but a smaller gap between optical remote sensing and SAR scenes, making the transfer to SAR more challenging. Validation on tasks like image-text retrieval shows the effectiveness of our progressive knowledge transfer strategy.
4.5.2 Effect of the Training Encoder
This ablation study examines the impact of image and text encoders on downstream tasks. Table 7 shows that the text encoder alone achieves poor performance, with 5.23% mean recall on SARVLM-1M and 1.49% accuracy on FUSAR. The image encoder alone improves performance to 9.57% mean recall and 99.62% accuracy on MSTAR-SOC. Combining both encoders yields the best results, with 15.25% mean recall on SARVLM-1M and 99.83% accuracy on SOC, while zero-shot accuracy on FUSAR rises to 12.41%. This confirms that cross modal contrastive learning is key for better generalization.
Figure 7: Visualization of results on semantic localization task from SARCLIP (ViT-L-14).
Figure 8: Visualization of SAR imagery caption results from SARCap.
4.5.3 Effect of the training layer
As shown in Figure 9, we evaluate the impact of varying the number of activated layers in the image and text encoders on the SARVLM-1M test set. The results reveal a consistent trend: model performance improves as more layers are activated. This highlights the importance of deep, expressive representations in both modalities for effective SAR image-text understanding.
4.6 Visualization
As shown in Figure 6, we use t-SNE [t-SNE] to visualize SARCLIP (ViT-L-14) features on three downstream datasets. The well-separated clusters indicate strong feature discrimination and meaningful semantic representations.
Figure 9: Ablation study on training layers of SARCLIP on the SARVLM-1M test set (Mean Recall %).
@@ 4.6 Visualization
Table 7: Ablation study result on several downstream tasks. Img. Enc. and Text Enc. denote the image encoder and text decoder respectively.(%)($\mathbf{\%}$)
Model Architecture SARVLM-1M test SOC VSA FUSAR
@@
5 Conclusion
We present SARVLM-1M, a large-scale SAR image-text dataset, and propose SARCLVLM, the first vision-language foundation model for SAR. Trained via a two-stage domain transfer strategy, SARCLIP effectively transfers knowledge from optical remote sensing data to SAR imagery, achieving strong multi-modal alignment, target recognition and zero-shot and semantic localization capability while SARCap performs well in SAR imagery captioning. In the future, we will explore integrating multi-modal large language models and agent techniques for SAR image interpretation.